NvidiaIntely (nvidiaintely):

The percent composition of sulfuric acid, H2SO4, is: 2% hydrogen, 33% sulfur, 65% oxygen 1% hydrogen, 33% sulfur, 34% oxygen 65% oxygen, 33% sulfur, 2% hydrogen Cannot be determined from the information given

OpenStudy (kevin):

Mr of H2SO4 = 98 H = 2 * 1 / 98 * 100% = 2.04% S = 1 * 32/ 98 * 100% = 32.65% O = 4 *16/98 *100% = 65.31%

OpenStudy (welshfella):

add up the atomic mass of 2 h's aone S and 4 O's. will give you value of 1 mole of sulphuric acid,

OpenStudy (welshfella):

then you can find the $age composition od each element

NvidiaIntely (nvidiaintely):

Okay...

NvidiaIntely (nvidiaintely):

So its option A right Kevin?

NvidiaIntely (nvidiaintely):

@welshfella @TheSmartOne @Photon336